# Convergent transformations of visual representation in brains and models

### v1

@@ 3 Discussion
Our study establishes a general, data-driven approach for mapping information processing routes in the human brain. By combining inter-subject representational alignment, model-based comparisons, and explainability techniques, we identified a large-scale, stimulus-driven cortical network that is conserved across individuals and robust to task and dataset variation. This network, recruited during natural scene perception, is organized into two primary streams: a medial-ventral pathway for visually grounded scene analysis, and a lateral-dorsal pathway specialized for social and biological content. This result provides a data-driven refinement of the canonical dual-stream model of vision, functionally supporting recent proposals for a “third visual pathway”30, 56 and demonstrating a complex functional architecture that is integrated into a whole-cortex network.
The convergence we observed between brain and model representations—though not mechanistically identical—suggests that deep networks provide a rich "bag-of-features" capable of capturing multiple stages of cortical processing. This nuanced view is necessary to explain our findings for the high-level hubs. The Ventral hub, with its cumulative integration of both visual and abstract information, showed alignment across a wide range of model layers. The LOTC hub, in contrast, showed a highly specialized alignment: it corresponded almost exclusively with the deepest semantic layers and only for those stimuli containing biological content. These results indicate that a simple one-to-one hierarchical model cannot account for the distinct computational strategies of the cortex, where recurrence and functional specialization contrast with the strictly sequential architecture of the models used for comparison60, 61, 62. However, the strong correspondence between the alignment found across brains and with models suggests that despite these mechanistic differences, both systems are sensitive to the same underlying stimulus structure, captured in regions where this stimulus-driven information is most functionally relevant.
The observation that brains and deep networks converge on similar representational geometries likely reflects shared constraints for efficiently encoding the structured statistics of the environment 5, 6, 63. The principle of efficient coding posits that the optimal way to compress sensory information is to accurately predict it; therefore, any system that learns to predict the world will necessarily develop compressed, efficient representations. The convergence we observe across brains and models likely arises because both systems, through different optimization processes (evolution and gradient descent), are finding similar solutions to this fundamental problem. This is increasingly recognized as a general principle in both neuroscience and artificial intelligence, with modern models converging on similar latent features across tasks and modalities 64, 65. Characterizing shared representational spaces thus provides a principled way to understand how both biological and artificial systems learn to extract meaningful information from their environments 4.
Methodologically, this work underscores the potential of population-scale, data-driven analyses to move beyond static anatomical maps, enabling systematic tracing of functional routes in the brain. No single method used here would have sufficed; it is the combination of identifying a shared space across subjects, using model hierarchies to probe its content, and deploying explainability techniques to interpret the results that provides a complete picture. With increasing access to high-quality, large-scale datasets, this integrated approach provides a path for hypothesis generation across sensory, cognitive, and cross-modal domains. The integration of explainability methods further allows for direct interpretation of what dimensions or features are routed through specific pathways, with implications for theory and for applications such as decoding and brain-computer interfaces.
Limitations remain, including the need for deeper mechanistic models linking representational geometry to neural computation and behavior, and for extending these methods to causal and cross-modal analyses. As recent work highlights, high RSA scores can arise from non-mechanistic similarities when using complex stimuli with inherent confounds 59. While our explainability analysis is a first step toward mitigating this, future work will require more precise methods to isolate the specific features driving alignment and to bridge the gap from representational similarity to true mechanistic understanding. Nonetheless, this study demonstrates that combining shared representational geometry across individuals with deep model features is a powerful framework for mapping and interpreting brain function at scale.
Our findings highlight the importance of tracing not just where, but how and what information is routed through the brain. This provides a principled, generalizable methodology for the field—moving from describing correspondence to generating more precise, mechanistically testable hypotheses about information processing in the brain.
@@ 4.1 Functional-MRI data
For the main analyses of this study, we used data from the Natural Scenes Dataset (NSD)23, a high-resolution 7T fMRI dataset comprising recordings from 8 participants performing a visual memory task. On each trial, participants viewed a naturalistic image for 3 seconds and indicated whether it was old or new. The stimulus set included a wide variety of natural scenes, spanning indoor and outdoor environments, static object views, and complex multi-element situations, including a broad range of social contexts. Each participant completed between 32 and 40 sessions, with 750 trials per session conducted over the course of a year. For participants who completed all 40 sessions, each image was presented three times, yielding a total of 30,000 trials per participant. Full experimental details are provided in Allen et al. (2022)23.
Analyses were based on the 1.0-mm volumetric preprocessing of the NSD dataset and the corresponding single-trial $\beta$-estimates (version 3) 23. We used GLM-denoised single-trial $\beta$-estimates 66, as their image-related information content has been validated in multiple decoding and image reconstruction studies 42, 67, 68. Voxels were grouped into cortical parcels using the HCP-MMP1.0 atlas 31, which defines 180 cortical regions per hemisphere. We employed the pre-aligned parcel masks provided by the NSD authors to extract voxel activity within each cortical region 23. Analyses were performed using both the symmetric version of the atlas (combining hemispheres) and the asymmetric version (preserving hemisphere separation); a supplementary analysis confirms that symmetric parcels display highly correlated representational profiles (Supplementary Fig. A1). In all figures, parcels are colored by macro-anatomical groups as defined in Huang et al. (2022) 47, to facilitate anatomical interpretation.
For each participant $p~{}\in~{}\mathcal{P}$, parcel $r~{}\in~{}\mathcal{R}$, and session $s~{}\in~{}\mathcal{T}p$, we define a response matrix $X_{p,r,s}~{}\in~{}\mathbb{R}^{n\times v_{p,r}}$ containing the voxel responses to each stimulus in that session. Here, $n$ denotes the number of trials (750 per session), and $v_{p,r}$ the number of voxels in parcel $r$ for participant $p$.
Comparative analyses were performed using fMRI data from the BOLD500024 and THINGS25 datasets. Preprocessing and analysis followed the same procedure described above, with minor adaptations to accommodate differences in stimulus sampling and repetition structure. Full details for these datasets are provided in Supplementary Section B.
All data analysed are fully de-identified and publicly available. The original NSD collection was approved by the University of Minnesota Institutional Review Board (Allen et al.23). The THINGS dataset was collected under NIH Institutional Review Board approval (Hebart et al.25), and BOLD5000 under the Institutional Review Board of Carnegie Mellon University (Chang et al.24). No new human-subjects data were acquired for this work.
@@ 4.2 Model Features Extraction
To compare how deep learning models encode visual and linguistic stimuli relative to fMRI responses, we selected vision and language models previously used by Huh et al. (2024) in studies of representational convergence across modalities 64. Model selection was conducted prior to analysis, balancing computational constraints and the goal of covering diverse training objectives and model scales. A full list of models is provided in Table 1.
We selected 17 vision models spanning families with distinct training regimes, including ViT AugReg69, CLIP70, 71, DINOv272, and MAE73. All architectures were based on sequential Vision Transformer (ViT) blocks74. Each NSD image was processed through the backbone of each model, and activations were extracted from the output of each ViT block (see Fig. 1A).
For language models, we selected 14 transformer-based encoders from the BLOOMZ75, Gemma 276, LLaMA77, 78, and LLaMA 379 families80. For each NSD image, we used text captions annotated by trained raters and originally sourced from the MS-COCO dataset81. Captions were processed through each model’s encoder, and activations were extracted at the output of each transformer block.
Feature extraction was performed using the Transformers82 and TIMM83 libraries, adapting procedures from Huh et al. (2024)64. All models were implemented in PyTorch84 and retrieved from the Hugging Face Model Hub.
@@
Vision ViT (AugReg)69 Supervised (21 K classes) ViT (AugReg) Tiny 10M 12 192
@@
CLIP (Vision)70, 71 Contrastive Image–Text CLIP (ViT) Base 86M 12 768
@@
Contrastive I-T + fine-tuning (12 K) CLIP (ViT) Base ft-12k 95M 12 768
CLIP (ViT) Large ft-12k 315M 24 1024
CLIP (ViT) Huge ft-12k 646 M 32 1280
DINOv272 Self-Supervised DinoV2 (ViT) Small 22M 12 384
@@
MAE73 Self-Supervised (Masked AE) MAE (ViT) Base 86M 12 768
@@
Language BLOOMZ75 Causal LM + Instruction FT BloomZ (560M) 559M 25 1024
@@
Gemma 276 Causal LM Gemma 2 (2B) 2.6B 27 2304
@@
LLaMA77, 78 Causal LM OpenLlama (3B) 3.4B 27 3200
@@
LLaMA 379 Causal LM Llama 3 (8B) 8B 33 4096
@@ 4.3 Alignment Measure
We used Representational Similarity Analysis (RSA)32 to quantify the correspondence between brain parcels and model layers geometries. For each system, the representational dissimilarity matrix (RDM) $D(X)$ was defined as $D(X)_{ij}=1-\rho(x_{i},x_{j})$, where $x_{i}$ and $x_{j}$ are rows of $X$ and $\rho$ is the Pearson correlation. The RSA score between two systems $X$ and $Y$ was the correlation between the upper triangles of their RDMs:
@@
Main analyses were replicated with Spearman correlation and Centered Kernel Alignment (CKA)85, yielding consistent results (see Supplementary Fig.A5). Details of the matrix reformulation that enabled efficient, large-scale RSA computations are provided in Supplementary Section C.
@@ 4.4 Inter-Subject alignment
Inter-subject RSA was computed using the subset of 1,000 images viewed by all NSD participants (the shared1000 set)23. Each image in this subset was presented up to three times per participant, with corresponding trial indices across sessions. Trials were uniquely identified by participant, image ID, and repetition index $k\in{1,2,3}$. For each pair of participants $p,q\in\mathcal{P}$ and cortical parcels $r,r^{\prime}\in\mathcal{R}$, we extracted the corresponding voxel responses $X_{p,r}$ and $X_{q,r^{\prime}}$ and computed RSA between matched trials.
Images shared across participants were always presented at the same trial positions for all participants, interleaved with unique images specific to each individual. To attenuate alignment driven by task or session structure rather than by true stimulus processing, we applied a cyclic repetition shift when matching trials across participants. Specifically, for each image $i$ and repetition index $k$, the trial $(i,k)$ in participant $p$ was matched with the trial $(i,(k\bmod 3)+1)$ in participant $q$, thereby preserving stimulus identity while disrupting repetition-locked confounds. This matching strategy and its impact are illustrated and compared in Supplementary Fig. A2. Let $\boldsymbol{i}_{pq}$ and $\boldsymbol{j}_{pq}$ denote the matched stimulus indices for participants $p$ and $q$, respectively. The inter-subject alignment between parcels $r$ and $r^{\prime}$ was then computed as:
@@
Group-level alignment was then obtained by averaging subject-level measures across participants. Cortical surface maps (e.g., Fig.2b) display group-level alignment, specifically comparing the same parcels across participant pairs ($r=r^{\prime}$). To benchmark the degree of shared geometry relative to individual consistency, we also compared inter-subject and within-subject alignment (see Supplementary Fig. D5), which yielded consistent results.
A parcel-by-parcel matrix of group-level inter-subject alignment values ($n_{\mathrm{rois}}\times n_{\mathrm{rois}}$) defined the whole-cortex connectivity network. We visualized the main network backbone and multiple major information routes using an iterative minimum spanning tree (k-MST) approach on the group-averaged connectivity matrix. At each iteration, a standard MST was computed (with edge weights defined as $1-\mathrm{similarity}$), penalizing previously selected edges to ensure the inclusion of new, critical pathways. The final backbone comprised the union of edges from all $k$ MSTs (see Supplementary Fig. D6), preserving multiple principal paths, avoiding arbitrary edge thresholding, and ensuring that all parcels remained connected while allowing for richer topology. MST computations were performed using the NetworkX and SciPy libraries86, 87.
Statistical significance was evaluated by comparing the observed alignment values to a null distribution generated by randomly shuffling stimulus labels (10,000 permutations, using a shared permutation scheme across all comparisons) with a two-tailed test. All p-values were FDR-corrected 88.

### v2

@@ 3 Discussion
In this study, we aimed to characterize how the human cortex encodes natural scenes by quantifying representational structure shared across individuals and comparing it to hierarchical features in DNNs. We identified a distributed network of cortical regions engaged by scene perception and characterized its major hubs and the stimulus features that organize the processing within it.
Two representational routes during scene viewing
We identified this distributed network engaged during scene perception using inter-subject representational connectivity, which captures similarity in how different regions organize the stimulus set. The resulting topology separated three prominent hubs connected by two dominant routes: an early visual hub, a ventral hub spanning ventromedial and parahippocampal scene regions, and an MT+/LOTC hub. The early visual hub bifurcated into branches toward the ventral and lateral hubs, and the ventral and lateral systems were additionally coupled via intermediate lateral occipital and ventral temporal parcels (Fig. 4). This organization does not match the classical ventral “what” versus dorsal “where/how” partition 35, 54, but is instead consistent with proposals that fractionate visual processing into additional routes, including a ventromedial pathway that emphasizes scene context and a dorsolateral pathway centered on MT+ and STS-adjacent regions that is associated with social/biological information 78, 70, 58.
The medial-ventral route aligns with accounts of a ventromedial “where/scene” pathway that supports environmental layout and spatial context, providing inputs to hippocampal circuitry77, 78, 76. In our data, parcels in this ventral hub showed broad correspondence to vision models across multiple depths (Fig. 3b), indicating that their shared geometry relates to features distributed across multiple stages of the vision-model hierarchy. This pattern held for scenes with and without biological agents (Fig. 5e–f). It was reduced in the object-centric THINGS dataset, which contains limited scene layout and context (Fig. 6e–f), confirming that this route is most detectable when the stimulus set contains rich environmental information. Language-model alignment in these regions showed no correspondence (Fig. 2d), diverging from findings using similarity-based models 94, 22, indicating that caption-derived similarity structure did not capture the visually grounded scene-layout geometry expressed along the medial-ventral route.
The lateral route aligns with accounts of an MT+/STS-centered pathway specialized for socially relevant information, including “third visual pathway” and dorsolateral subdivision proposals 78, 70, 58. Here, shared geometry was dominated by biological-agent content. The leading LOTC dimension separated scenes with animate agents from those without, and restricting analyses to images without biological agents reduced LOTC inter-subject alignment and weakened the lateral branch of the network (Fig. 5e-f; Supplementary Fig. 8). LOTC parcels aligned most strongly with later stages of vision models (Fig. 3c) and also aligned with language-model representations (Fig. 2d,f,g), consistent with this route emphasizing socially structured distinctions that are expressed in deep visual features and are also present in text-derived semantic structure.
Model-brain alignment across the visual hierarchy
Model-brain RSA related the representational geometry measured in cortex to the ordered progression of features across model depth. For vision models, model-brain correspondence tracked the inter-subject shared geometry across parcels and scaled with IS-RSA with an approximate power-law relationship (Fig. 2e). This suggests that the geometry shared across observers during scene viewing is also present in image-trained feature spaces, consistent with the view that a reduced set of coarse dimensions is shared between brains and models and accounts for a large fraction of brain-model correspondence in visual tasks 13.
Layer-wise profiles revealed a parallel with the posterior-to-anterior ordering expected from the visual hierarchy (Fig. 3) 20, 28, 54, but the mapping was not one-to-one. For example, while the ventral hub showed correspondence across multiple depths, the MT+/LOTC hub showed selective alignment restricted to specific stimulus categories. This is consistent with cortical representations drawing on features distributed across multiple stages and being modulated by stimulus content, rather than following a strictly feedforward depth-to-stage mapping of models.
Language models exhibited a distinct profile, reflecting the nature of their input. Because these models processed tokenized captions–abstracted from pixel data–they did not follow the cortical depth gradient observed in vision models. Instead, alignment was concentrated in LOTC and adjacent temporal regions, driven primarily by biological-agent content. Layer-wise alignment rose early and plateaued, scaling with tokenizer alignment (Supplementary Note 5). This pattern suggests a correspondence driven by coarse semantic content–closer to a “bag-of-concepts” organization than to progressively composed syntax 100–indicating convergence on LOTC dimensions that are independent of low-level visual features.
More generally, these findings align with observations that modern models tend to converge on similar latent representations across objectives and architectures as they scale 13, 49, 50. One interpretation is that shared stimulus statistics, combined with efficiency constraints 31, favor the emergence of common representational axes that organize natural inputs by capturing structure essential for prediction and compression.
Reliability and interpreting RSA magnitudes
Across analyses, vision-model alignment closely tracked inter-subject alignment (Fig. 2e). RSA variants that differ in their noise sensitivity (e.g., within-subject RSA or alternative repetition-matching schemes) were also related by approximate power-law scaling (Supplementary Note 4). This pattern is consistent with these measures indexing a common stimulus-locked representational geometry, differing primarily in how strongly measurement noise attenuates the observed correlations.
This explains cases where model-brain RSA numerically exceeded IS-RSA in parcels with low-to-moderate reliability (e.g., posterior cingulate cortex; Supplementary Fig. 1). IS-RSA correlates two noisy brain estimates, whereas model-brain RSA pairs one noisy estimate with a effectively noiseless model RDM. When reliability is limited, the noise from both sources in IS-RSA compounds the attenuation, lowering the observed correlation more than in model-brain RSA, even if both track the same underlying geometry.
These results indicate that raw IS-RSA values may not always serve as an appropriate numeric ceiling for model-brain correspondence 63, 21, 92. Instead, RSA magnitudes are most informative when interpreted relative to reliability. Crucially, while adequate signal quality is a prerequisite for observing shared geometry–meaning the spatial extent of these maps inherently reflects the distribution of reliable fMRI signal–reliability alone cannot explain the topological organization we observed. The recovery of dissociable streams with distinct feature selectivities (e.g., ventral vs. lateral) confirms that this framework identifies specific information-processing networks, rather than merely tracking cortex-wide variations in signal quality. The empirical scaling relations provided here offer a practical calibration for separating these effects.
Conclusions and scope
In this work, we characterized a cortical network engaged during scene viewing at the level of representational geometry. By integrating complementary analyses of alignment, representational connectivity, and dimensionality decomposition, we linked three questions often treated separately: where shared stimulus information is expressed, how that information is routed across regions, and which stimulus properties organize representations within different network hubs. Together, these analyses provide a comprehensive perspective on how scene information is structured across the human cortex.
This representational approach offers a general and adaptable methodology for identifying task-engaged networks and comparing information transformation across diverse systems. However, this flexibility entails a tradeoff: inferences concern representational organization rather than the precise mechanisms that implement it 3, 9. Because these comparisons rely on global similarity structure, computational dimensions that are subtle but behaviorally relevant may contribute weakly to the measured geometry, or be missed entirely if not captured by the chosen metric or feature sampling 57. Consequently, these results should be interpreted as context-dependent–varying with stimulus sampling, task demands, and measurement reliability 25–motivating future work to causally isolate the specific stimulus factors driving these alignments.
Despite these limitations, this framework provides a practical route for leveraging large-scale data to move beyond simply asking whether two systems encode information similarly. By decomposing shared geometry to isolate driving features, the approach helps identify candidate factors and interactions that explain what information is shared across cortex and how it is distributed. As dataset scale and reliability increase, these tools can bridge the gap between network-level descriptions and experiments probing how distributed cortical systems support perception and flexible behavior in dynamic environments.
@@ 4.1 Functional-MRI data
For the main analyses, we used the Natural Scenes Dataset (NSD) 2, a high-resolution 7T fMRI dataset with 8 participants performing an image recognition-memory task. On each trial, participants viewed a natural scene for 3s and reported whether it was old or new. Each participant completed 32–40 scanning sessions (750 trials per session), yielding up to $\sim 30,000$ trials per participant with repeated presentations of images across sessions. Full acquisition and experimental details are reported in Allen et al. 2.
Analyses were based on the 1.0-mm volumetric preprocessing of the NSD dataset and the corresponding single-trial $\beta$-estimates (version 3) 2. We used GLM-denoised single-trial $\beta$-estimates 73, as their image-related information content has been validated in multiple decoding and image reconstruction studies 87, 65, 81. Voxels were grouped into cortical parcels using the HCP-MMP1.0 atlas 34, which defines 180 cortical regions per hemisphere. We employed the pre-aligned parcel masks provided by the NSD authors to extract voxel activity within each cortical region 2. Analyses were performed using both the symmetric version of the atlas (combining hemispheres) and the asymmetric version (preserving hemisphere separation); a supplementary analysis confirms that symmetric parcels display highly correlated representational profiles (Supplementary Note 3). In all figures, parcels are colored by macro-anatomical groups as defined in Huang et al. (2022) 47, to facilitate anatomical interpretation.
[…]
Comparative analyses were performed using fMRI data from the BOLD5000 11, and THINGS 46 datasets. We selected these datasets because IS-RSA and KMCCA require a large number of matched stimuli across participants, which is often limited in larger-cohort naturalistic datasets. Preprocessing and analysis followed the same procedure described above, with minor adaptations to accommodate differences in stimulus sampling and repetition structure. Details are provided in Supplementary Note 8.
All data analysed are fully de-identified and publicly available. The original NSD collection was approved by the University of Minnesota Institutional Review Board (Allen et al.2). The THINGS dataset was collected under NIH Institutional Review Board approval (Hebart et al.46), and BOLD5000 under the Institutional Review Board of Carnegie Mellon University (Chang et al.11). No new human-subjects data were acquired for this work.
@@ 4.2 Model Features Extraction
To compare how deep learning models encode visual and linguistic stimuli relative to fMRI responses, we selected vision and language models previously used by Huh et al. (2024) in studies of representational convergence across modalities 49. Model selection was conducted prior to analysis, balancing computational constraints and the goal of covering diverse training objectives and model scales. A full list of models is provided in Table 1.
We selected 17 vision models spanning families with distinct training regimes, including ViT AugReg84, CLIP74, 14, DINOv264, and MAE44. All architectures were based on sequential Vision Transformer (ViT) blocks24. Each NSD image was processed through the backbone of each model, and activations were extracted from the output of each ViT block (see Fig. 1A).
For language models, we selected 14 transformer-based encoders from the BLOOMZ61, Gemma 232, LLaMA88, 33, and LLaMA 336 families89. For each NSD image, we used text captions annotated by trained raters and originally sourced from the MS-COCO dataset59. Captions were processed through each model’s encoder, and activations were extracted at the output of each transformer block.
Feature extraction was performed using the Transformers97 and TIMM96 libraries, adapting procedures from Huh et al.49. All models were implemented in PyTorch67 and retrieved from the Hugging Face Model Hub.
@@
Vision ViT (AugReg)84 Supervised (21 K classes) ViT (AugReg) Tiny 10M 12 192
@@
CLIP (Vision)74, 14 Contrastive Image–Text CLIP (ViT) Base 86M 12 768
@@
Contrastive I-T + fine‑tuning (12 K) CLIP (ViT) Base ft‑12k 95M 12 768
CLIP (ViT) Large ft‑12k 315M 24 1024
CLIP (ViT) Huge ft‑12k 646 M 32 1280
DINOv264 Self‑Supervised DinoV2 (ViT) Small 22M 12 384
@@
MAE44 Self‑Supervised (Masked AE) MAE (ViT) Base 86M 12 768
@@
Language BLOOMZ61 Causal LM + Instruction FT BloomZ (560M) 559M 25 1024
@@
Gemma 232 Causal LM Gemma 2 (2B) 2.6B 27 2304
@@
LLaMA88, 33 Causal LM OpenLlama (3B) 3.4B 27 3200
@@
LLaMA 336 Causal LM Llama 3 (8B) 8B 33 4096
@@ 4.3 Alignment Measure
We used Representational Similarity Analysis (RSA)55 to quantify the correspondence between brain parcels and model layers geometries. For each system, the representational dissimilarity matrix (RDM) $D(X)$ was defined as $D(X)_{ij}=1-\rho(x_{i},x_{j})$, where $x_{i}$ and $x_{j}$ are rows of $X$ and $\rho$ is the Pearson correlation. The RSA score between two systems $X$ and $Y$ was the correlation between the upper triangles of their RDMs:
@@
Main analyses were replicated with Spearman correlation and Centered Kernel Alignment (CKA)53, yielding consistent results (see Supplementary Note 6). Details of the matrix reformulation that enabled efficient, large-scale RSA computations are provided in Supplementary Note 9.
@@ 4.4 Inter-Subject alignment
Inter-subject RSA was computed using the subset of 1,000 images viewed by all NSD participants (the shared1000 set)2. Each image in this subset was presented up to three times per participant, with corresponding trial indices across sessions. Trials were uniquely identified by participant, image ID, and repetition index $k\in{1,2,3}$. For each pair of participants $p,q\in\mathcal{P}$ and cortical parcels $r,r^{\prime}\in\mathcal{R}$, we extracted the corresponding voxel responses $X_{p,r}$ and $X_{q,r^{\prime}}$ and computed RSA between matched trials.
Images shared across participants were always presented at the same trial positions for all participants, interleaved with unique images specific to each individual. To attenuate alignment driven by task or session structure rather than by stimulus processing, we applied a cyclic repetition shift when matching trials across participants. Specifically, for each image $i$ and repetition index $k$, the trial $(i,k)$ in participant $p$ was matched with the trial $(i,(k\bmod 3)+1)$ in participant $q$, thereby preserving stimulus identity while disrupting repetition-locked confounds. This matching strategy and its impact are illustrated and compared in Supplementary Note 1. Let $\boldsymbol{i}_{pq}$ and $\boldsymbol{j}_{pq}$ denote the matched stimulus indices for participants $p$ and $q$, respectively. The inter-subject alignment between parcels $r$ and $r^{\prime}$ was computed as:
@@
Group-level alignment was then obtained by averaging subject-level measures across participants. Cortical surface maps (e.g., Fig.2b) display group-level alignment, specifically comparing the same parcels across participant pairs ($r=r^{\prime}$). To benchmark the degree of shared geometry relative to individual consistency, we also compared inter-subject and within-subject alignment (see Supplementary Fig. 3), which yielded consistent results.
A parcel-by-parcel matrix of group-level inter-subject alignment values ($n_{\mathrm{rois}}\times n_{\mathrm{rois}}$) defined the whole-cortex connectivity network. We visualized the main network backbone and multiple major representation routes using an iterative minimum spanning tree (k-MST) approach on the group-averaged connectivity matrix. At each iteration, a standard MST was computed (with edge weights defined as $1-\mathrm{similarity}$), removing previously selected edges to ensure the inclusion of new pathways. The final backbone comprised the union of edges from all $k$ MSTs (see Supplementary Fig. 7), preserving multiple principal paths, avoiding arbitrary edge thresholding, and ensuring that all parcels remained connected while allowing for richer topology. MST computations were performed using the NetworkX and SciPy libraries91, 39.
Statistical significance was evaluated by comparing the observed alignment values to a null distribution generated by randomly shuffling stimulus labels (10,000 permutations, using a shared permutation scheme across all comparisons) with a two-tailed test. All p-values were FDR-corrected 7.